# On Three-Layer Data Markets

### v2

@@ 1 Introduction
With the advent of new functionalities by information technology companies such as Apple or Google that reduce or eliminate tracking via third-party cookies, there has been an increased reliance on first-party data, which is information gathered by a company from its own user base. For example, as elaborated in Cross [2023b], major retailers such as BestBuy and Walmart capitalize on this data through loyalty programs, customer purchase records, and subscriptions. They do this by selling the data to advertisers who aim to place targeted ads on the retailers’ websites or across various online platforms. In a similar vein, Cross [2023a] notes that Mastercard, a leading payment technology firm, trades cardholder transaction information via third-party online data marketplaces and its internal “Data and Services” division. This practice grants numerous entities access to extensive consumer data and insights. For instance, their Intelligent Targeting service allows businesses to leverage “Mastercard ${}^{\circ}$ data insights” to create and execute advertising campaigns that specifically target potential “high-value” customers.
In this complex digital ecosystem, users face the dilemma of benefiting from the services provided by these platforms while potentially compromising their privacy. Platforms, on the other hand, must balance the profitability derived from data sharing with the need to maintain user trust and satisfaction. This interplay between users’ privacy concerns and the economic incentives of platforms raises important questions about the sustainability of current data market practices. Notable regulatory actions have been taken to improve user experiences. For instance, the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA) have been introduced to protect user privacy and to limit the amount of data shared with the platforms. Although, at first sight, this helps users, there have been several studies pointing to its detrimental impact both on platforms and users. This is because these regulations have led to a more concentrated market structure by limiting competition in data markets. This results from the regulation’s restrictions on data sharing, which may slow down data-based innovations and services. For instance, studies have shown that GDPR has led to a significant reduction in the number of available apps in the market, with new app entries falling by half following the regulation’s implementation (Janssen et al. [2022]). This decline in market entries indicates a potential loss of innovation as new and smaller players find it increasingly difficult to compete under these regulations. This effect is particularly detrimental to small companies that rely on these data synergies for innovation and market growth (see also Gal and Aviv [2020]).
@@ 1.1 Our model and main results
We model the interaction between the players as a multi-stage game in which, in the first stage, the platforms decide whether they want to provide services to the user and the privacy level they want to deliver to the data buyer (so that the user is incentivized to share). We model the privacy level that the platform provides as a noise level added to the user data. Adding a suitable level of noise is common practice in, for example, the analysis of census data Abowd [2018] and is theoretically justified in the literature on differential privacy Dwork et al. [2014].
@@
We adopt the subgame Nash equilibrium as our equilibrium concept and use backward induction to characterize it. Finding the equilibrium of this game is complex mainly because the platforms’ choice of noise level not only directly impacts the user’s utility but also changes the data buyer’s decision, which indirectly impacts the user’s utility. Also, the competition among platforms further complicates the dynamic of the game. For instance, a platform may still be incentivized to sell data at a noise level higher than needed for the user to be willing to share their data. At first glance, this may seem unreasonable as it reduces the quality of the data that they can sell, and hence, the data buyer would pay less for it. However, increasing the privacy could persuade the user to forgo other platforms that offer less stringent privacy guarantees and only use this platform’s service. As a result, this can give a monopoly to the platform in terms of access to the user’s data, and hence, the data buyer would end up paying more to them.
@@
We demonstrate that these findings extend to the general scenario with $K>2$ platforms. Specifically, we establish that when $\beta$ exceeds a certain threshold, signifying the data buyer’s willingness to pay more for the user data, all platforms will enter the market and provides services to the user. Conversely, when $\beta$ is below this threshold, only low-cost platforms will participate, while high-cost ones will abstain. Although competition does not favor the user, it results in positive utility for the data buyer. Moreover, the overall utilitarian welfare increases proportionally with the number of participating platforms.
@@
Building on our equilibrium characterization, we then study regulations that aim to improve user utility. Here, our analysis provides three main insights. Firstly, we consider a ban on platforms sharing user data with the data buyers. This regulation only maximizes the user utility if all the platforms are low-cost. Specifically, under this regulation, users obtain the best of both worlds: all platforms (that have a low cost for providing services) enter the market, and the user benefits from the services of all of them while incurring zero privacy loss. This observation suggests that if we only had large firms whose costs of providing service per user are small, then the draconian regulation that involves banning data sharing would have been optimal. However, this is no longer the case if the market includes both high and low-cost platforms.
In this case, our second insight reveals that imposing a uniform minimum privacy mandate improves user utility, particularly when the value of the user data to the data buyer is significant enough for platforms to charge a high price for it. This means that all platforms need to perturb the user data by adding to it a noise with a high enough noise level. Here, our analysis identifies specific conditions where GDPR-type policies, which limit user data usage, can benefit users. Finally, our third insight suggests that, despite potential implementation challenges, a non-uniform minimum privacy mandate — banning data sharing for low-cost platforms while imposing a minimum privacy standard for high-cost platforms — further enhances user utility. This regulation entails prohibiting large, low-cost platforms from sharing user data, while permitting smaller, high-cost platforms to do so, albeit in a limited manner through privacy mechanisms. This approach aligns with our earlier discussion that GDPR-type regulations disproportionately harm small businesses. Therefore, a non-uniform regulation favoring small businesses is preferred not only by these businesses but also by the users.
@@ 1.2 Related literature
Our paper broadly relates to the emerging literature on data markets and online platform behavior. Earlier works on this topic include Acquisti et al. [2016], Posner and Weyl [2018], Ali et al. [2019], Jones and Tonetti [2020], and Dosis and Sand-Zantman [2022] which study the consequences of protecting and disclosing personal information about individuals. Additionally, Acemoglu et al. [2022], Bergemann et al. [2020], Ichihashi [2020b], and Fainmesser et al. [2022] study the privacy consequences of data externality (whereby a user’s data reveals information about others). Acemoglu et al. [2023a] study user-optimal privacy-preserving mechanisms that platforms can adopt to balance privacy and learning and Acemoglu et al. [2023b] develop an experimentation game to study the implications of the platform’s information advantage in product offering. Finally, Ichihashi [2021] studies data intermediaries, and Ichihashi [2020a] studies consumer privacy choices while online platforms adjust their privacy guarantees dynamically to incentivize more data sharing from consumers.
@@
More closely related to ours are papers that study data sharing among users, platforms, and third-party data buyers and, in particular, the impacts of banning such data sharing. In this regard, Bimpikis et al. [2021] and Argenziano and Bonatti [2023] develop a two-round model of the interaction between a user and two competing platforms that can collect data on user preferences (and potentially use it for price discrimination). Bimpikis et al. [2021] show that banning data sharing can hurt the user when the platforms offer complementary products, and Argenziano and Bonatti [2023] show that banning data sharing can hurt the user if the benefits of knowing user preferences by the seller and personalized services are limited. The above papers consider the interactions between a user and a platform that itself uses the user data for price discrimination and, therefore, can potentially hurt the user. In contrast, we study a different problem that relates to how platforms sell user data to other third-party data buyers. Therefore, in addition to the differences in modeling choices and analysis, we depart from the above papers by considering a three-layer data market that includes the interaction between the platforms and the data buyer and focusing on deriving insights on effective regulations.
@@
Also related to our work is Ravichandran and Korula [2019] who empirically study the effect of the data-sharing ban on platforms’ revenues and Madsen and Vellodi [2023] who study the effects of a complete ban on innovation (we refer to Pino [2022] for a comprehensive survey on the microeconomics of data). We depart from these papers by focusing on the impact of regulations (and in particular data-sharing ban) on user’s utility. Our paper also relates to the literature that studies various forms of monetizing user data by platforms, such as selling cookies Bergemann and Bonatti [2015], efficient pricing for large data sets Abolhassani et al. [2017], dynamic sales Immorlica et al. [2021], Drakopoulos and Makhdoumi [2023], monetization while ensuring the data is replicable Falconer et al. [2023], and the design and price of information where a data buyer faces a decision problem under uncertainty and can augment their initial private information with supplemental data from a data seller Bergemann et al. [2018].
@@
Finally, our paper relates to the growing literature on data acquisition and the design of mechanisms for incentivizing data sharing, including Ghosh and Roth [2011], Ligett and Roth [2012], Nissim et al. [2014], Cummings et al. [2015], Chen et al. [2018], Chen and Zheng [2019], Fallah et al. [2023], Cummings et al. [2023], Fallah et al. [2022], and Karimireddy et al. [2022]. We depart from this line of work as we consider a three-layer market where the privacy cost occurs due to selling data to a third-party buyer.
@@
The rest of the paper proceeds as follows. In Section 2, we describe our model and the interactions among the platforms, the data buyer, and the user. We also present some properties of the revealed information measure in our setting such as its monotonicity and submodularity. In Section 3, we introduce our equilibrium concept and provide preliminary characterizations of it. In Section 4, we focus on a setting with two platforms and provide a full characterization of the equilibrium. The focus on two platforms allows us to describe some of the nuances of the interactions in our model. In Section 5, we establish that our main characterizations continue to hold in the general setting with any number of platforms. We also provide several insights and comparative statics, leading to our discussion of the regulations in Section 6. Finally, Section 7 concludes while the appendix contains all the omitted proofs from the text.
@@ 2 Model
We consider $K\in\mathbb{N}$ platforms that are interacting with a user and a data buyer. The user benefits from using the services of these platforms. In return, the platforms acquire data about the user’s preferences and behavior. The platforms may subsequently sell this data to a third party, such as an advertiser, which we refer to as a data buyer (or, interchangeably, a buyer). The data buyer is interested in learning about the user’s preferences, which can adversely impact the user’s utility. This negative impact can be because of price discrimination, unfairly targeted advertising, manipulation, or intrinsic privacy losses. In the example of MasterCard described in the introduction, MasterCard is one of the platforms, and the data buyer is any other firm that purchases user data from MasterCard.
Formally, we represent the user’s data by $\bm{\theta}\in\mathbb{R}^{d}$, which follows a Gaussian distribution with zero mean and identity variance, that is, $\bm{\theta}\sim\mathcal{N}(0,\mathbb{I}_{d})$. This vector $\bm{\theta}$ can be interpreted as the user’s feature vector. Each platform $i\in[K]\triangleq\{1,\dots,K\}$ offers a service whose quality depends on how accurately the platform can learn the match between the user’s feature vector and the platform’s characteristics, i.e.,
$\bm{x}_{i}^{\top}\bm{\theta}.$
Here, $\bm{x}_{i}\in\mathbb{R}^{d}$ is a known unit-norm vector symbolizing the characteristic vector of platform $i$.
@@
The introduced noise ensures a degree of privacy concerning the user’s data. [^1] We set the noise variance to one to simplify the results. However, the analysis remains valid for any (potentially heterogeneous) level of noise.
The data buyer aims to estimate a linear function of the user’s data, $\bm{y}^{\top}\bm{\theta}$. Here, $\bm{y}\in\mathbb{R}^{d}$ with $\|\bm{y}\|=1$ denotes the data buyer’s known characteristic vector. If platform $i$ possesses the user’s data (i.e., $a_{i}=1$), they present the data buyer with an option to acquire a noisy version of their data at a price $p_{i}$. Therefore, the data buyer strategy profile is $\bm{b}\in\{0,1\}^{K}$, with $b_{i}=1$ denoting their acceptance of the platform $i$’s offer. The data buyer then receives a signal
@@
for the price $p_{i}$, where $\sigma_{i}^{2}$ represents the variance of the noise that platform $i$ adds to the data they share with the data buyer.
@@
We set $\tilde{s}_{i}=\text{NA}$ when the data buyer does not obtain any information from platform $i$. This can occur either because the user has not shared their data with platform $i$ or because the data buyer declines the offer from platform $i$ or the platform has not provided the service to the user. In this case, $p_{i}$ is irrelevant, and we set it to zero.
@@
denote the correlation between the user $i$’s data and the data buyer. Then, we assume
@@
Note that under this assumption, the data offered by different platforms are viewed as substitutes for each other from the data buyer’s perspective. However, if this assumption does not hold, a complementary effect could also be observed. To illustrate this, suppose we have $K=2$ platforms where $x_{1}$ is orthogonal to $y$, i.e., $\gamma_{1}=0$, and also $x_{2}=\gamma_{2}y+(1-\gamma_{2})x_{1}$. In this scenario, the data from platform 1 alone has no value to the data buyer, as $x_{1}^{\top}\theta$ provides no information about $\theta^{\top}y$. However, if the data buyer decides to purchase data from the second platform, then the data from the first platform could also become valuable. To understand why, observe that the data from the second platform is a convex combination of two components: $x_{1}^{\top}\theta$ and $y^{\top}\theta$. Therefore, if the data buyer acquires the data from the second platform, learning $x_{1}^{\top}\theta$ could assist in better distinguishing these two components. In other words, the data from the first platform is valuable only if accompanied by the data from the second platform. This scenario exemplifies the complementary effect we described earlier. Assumption 1 excludes this possibility, allowing us to primarily focus on the substitution effect. This assumption also implies that the correlation among data for different platforms is given by
$\mathbb{E}[\bm{x}_{i}\bm{x}_{j}]=\gamma_{i}\gamma_{j}.$

### v3

@@ 1 Introduction
With the advent of new functionalities by information technology companies such as Apple or Google that reduce or eliminate tracking via third-party cookies, there has been an increased reliance on first-party data, which is information gathered by a company from its own user base. For example, as elaborated by Cross [2023b], major retailers such as BestBuy and Walmart capitalize on this data through loyalty programs, customer purchase records, and subscriptions. They do this by selling the data to advertisers who aim to place targeted ads on the retailers’ websites or across various online platforms. In a similar vein, Cross [2023a] notes that Mastercard, a leading payment technology firm, trades cardholder transaction information via third-party online data marketplaces and its internal “Data and Services” division. This practice grants numerous entities access to extensive consumer data and insights. For instance, their Intelligent Targeting service allows businesses to leverage “Mastercard 360 data insights” to create and execute advertising campaigns that specifically target potential “high-value” customers.
In this complex digital ecosystem, users face the dilemma of benefiting from the services provided by these platforms while potentially compromising their privacy. Platforms, on the other hand, must balance the profitability derived from data sharing with the need to maintain user trust and satisfaction. This interplay between users’ privacy concerns and the economic incentives of platforms raises important questions about the sustainability of current data market practices. Notable regulatory actions have been taken to improve user experiences. For instance, the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA) have been introduced to protect user privacy and to limit the amount of data shared with the platforms. Although, at first sight, this helps users, there have been several studies pointing to its detrimental impact both on platforms and users. This is because these regulations have led to a more concentrated market structure by limiting competition in data markets. This results from the regulation’s restrictions on data sharing, which may slow down data-based innovations and services. For instance, studies have shown that GDPR has led to a significant reduction in the number of available apps in the market, with new app entries falling by half following the regulation’s implementation [see, e.g., Aridor et al., 2021, Janssen et al., 2022]. This decline in market entries indicates a potential loss of innovation as new and smaller players find it increasingly difficult to compete under these regulations. This effect is particularly detrimental to small companies that rely on these data synergies for innovation and market growth [Gal and Aviv, 2020]. On a similar note, Demirer et al. [2024] also show that GDPR, on average, represents a 20% increase in the cost of data, with the “distortionary effects of the GDPR are highest for the smallest firms.”
@@ 1.1 Our model and main results
We model the interaction between the players as a multi-stage game in which, in the first stage, the platforms decide whether they want to provide services to the user and the privacy level they want to deliver to the data buyer (so that the user is incentivized to share). We model the privacy level that the platform provides as a noise level added to the user data. Adding a suitable level of noise is the common practice for achieving privacy, primarily motivated by the differential privacy framework of Dwork et al. [2014]. This approach is widely used both in the machine learning literature and in real-world applications, ranging from census data [see, e.g., Abowd, 2018] to technology companies such as Google and Apple [see, e.g., Erlingsson et al., 2014, differential privacy team, 2017].
@@
We adopt the subgame perfect equilibrium as our equilibrium concept and use backward induction to characterize it. Finding the equilibrium of this game is complex mainly because the platforms’ choice of noise level not only directly impacts the user’s utility but also changes the data buyer’s decision, which indirectly impacts the user’s utility. Also, the competition among platforms further complicates the dynamic of the game. For instance, a platform may still be incentivized to sell data at a noise level higher than needed for the user to be willing to share their data. At first glance, this may seem unreasonable as it reduces the quality of the data that they can sell, and hence, the data buyer would pay less for it. However, increasing the privacy could persuade the user to forgo other platforms that offer less stringent privacy guarantees and only use this platform’s service. As a result, this can give a monopoly to the platform in terms of access to the user’s data, and hence, the data buyer would end up paying more to them.
@@
We demonstrate that these findings extend to the general scenario with $K>2$ platforms. Specifically, we establish that when $\beta$ exceeds a certain threshold, signifying the data buyer’s willingness to pay more for the user data, all platforms will enter the market and provide services to the user. Conversely, only low-cost platforms will participate when $\beta$ is below this threshold, while high-cost platforms will abstain. Although competition does not favor the user, it results in positive utility for the data buyer. Moreover, the overall utilitarian welfare increases proportionally with the number of participating platforms.
@@
Building on our equilibrium characterization, we then study regulations that aim to improve user utility. Here, our analysis provides three main insights. Firstly, we will consider a ban on platforms sharing user data with data buyers. This regulation only maximizes the user utility if all the platforms are low-cost. Specifically, under this regulation, users obtain the best of both worlds: all platforms (that have a low cost for providing services) enter the market, and the user benefits from the services of all of them while incurring zero privacy loss. This observation suggests that if we only had large firms whose costs of providing service per user are small, then the draconian regulation that involves banning data sharing would have been optimal. However, this is no longer the case if the market includes both high- and low-cost platforms.
In this case, our second insight reveals that imposing a uniform minimum privacy mandate improves user utility, particularly when the value of the user data to the data buyer is significant enough for platforms to charge a high price for it. This means that all platforms need to perturb the user data by adding to it a noise with a high enough noise level. Here, our analysis identifies specific conditions in which GDPR-type policies, which limit user data usage, can benefit users. Finally, our third insight suggests that, despite potential implementation challenges, a non-uniform minimum privacy mandate—banning data sharing for low-cost platforms while imposing a minimum privacy standard for high-cost platforms—further enhances user utility. This regulation prohibits large, low-cost platforms from sharing user data while permitting smaller, high-cost platforms to do so, albeit in a limited manner through privacy mechanisms. This approach aligns with our earlier discussion that GDPR-type regulations disproportionately harm small businesses. Therefore, a non-uniform regulation favoring small businesses is preferred not only by these businesses but also by the users.
@@ 1.2 Related literature
Our paper broadly relates to the emerging literature on data markets and online platform behavior. Earlier works on this topic include Acquisti et al. [2016], Posner and Weyl [2018], Ali et al. [2019], Jones and Tonetti [2020], and Dosis and Sand-Zantman [2022], all of whom study the consequences of protecting and disclosing personal information about individuals. Additionally, Acemoglu et al. [2022], Bergemann et al. [2020], Ichihashi [2020b], and Fainmesser et al. [2022] study the privacy consequences of data externality (whereby a user’s data reveals information about others). Acemoglu et al. [2023a] study user-optimal privacy-preserving mechanisms that platforms can adopt to balance privacy and learning and Acemoglu et al. [2023b] develop an experimentation game to study the implications of the platform’s information advantage in product offering. Finally, Ichihashi [2021] studies data intermediaries, and Ichihashi [2020a] studies consumer privacy choices while online platforms dynamically adjust their privacy guarantees to incentivize more consumer data sharing.
@@
More closely related to ours are papers that study data sharing among users, platforms, and third-party data buyers and, in particular, the impacts of banning such data sharing. In this regard, Bimpikis et al. [2021] and Argenziano and Bonatti [2023] develop a two-round model of the interaction between a user and two competing platforms that can collect data on user preferences (and potentially use it for price discrimination). Bimpikis et al. [2021] show that banning data sharing can hurt the user when the platforms offer complementary products, and Argenziano and Bonatti [2023] show that banning data sharing can hurt the user if the benefits of knowing user preferences by the seller and personalized services are limited. The above papers consider the interactions between a user and a platform that itself uses the user data for price discrimination and, therefore, can potentially hurt the user. In contrast, we study a different problem that relates to how platforms sell user data to other third-party data buyers. Therefore, in addition to the differences in modeling choices and analysis, we depart from the above papers by considering a three-layer data market that includes the interaction between the platforms and the data buyer and focusing on deriving insights on effective regulations. Furthermore, Fang and Kim [2021] and Bergemann and Bonatti [2024] develop a model to analyze the impact of competition on how data-rich platforms can extract value from sellers through advertising.
@@
Also related to our work is Ravichandran and Korula [2019], who empirically study the effect of the data-sharing ban on platforms’ revenues, and Madsen and Vellodi [2023], who study the effects of a complete ban on innovation (we refer to Pino [2022] for a comprehensive survey on the microeconomics of data). We depart from these papers by focusing on the impact of regulations (and in particular, data-sharing ban) on users’ utility. Our paper also relates to the literature that studies various forms of monetizing user data by platforms, such as selling cookies [Bergemann and Bonatti, 2015], efficient pricing for large data sets [Abolhassani et al., 2017], dynamic sales [Immorlica et al., 2021], [Drakopoulos and Makhdoumi, 2023], monetization while ensuring the data is replicable [Falconer et al., 2023], and the design and price of information where a data buyer faces a decision problem under uncertainty and can augment their initial private information with supplemental data from a data seller [Bergemann et al., 2018].
@@
More broadly, our paper relates to the growing literature on data acquisition and the design of mechanisms for incentivizing data sharing [Ghosh and Roth, 2011, Ligett and Roth, 2012, Nissim et al., 2014, Cummings et al., 2015, Chen et al., 2018, Chen and Zheng, 2019, Fallah et al., 2023, Cummings et al., 2023, Fallah et al., 2022, Karimireddy et al., 2022]. We depart from this line of work as we consider a three-layer market where the privacy cost occurs due to selling data to a third-party buyer.
Finally, our paper relates to the supply chain models with competition in the intermediate layer. In these models, upstream and downstream firms interact through contractual arrangements, and the middle layer competes either in prices or quantities (see, e.g., Section 5 of [Cachon, 2003]). Related work by Lariviere [1999], Corbett et al. [2004], and Bernstein and Federgruen [2005] explores how such competition shapes equilibrium prices, efficiency, and welfare. In these settings, competition among intermediaries tends to benefit the downstream buyer and can reduce the intermediaries’ profits—an observation that parallels our finding that platform competition benefits the data buyer but not the users. Despite this structural analogy, our framework departs from the supply chain literature in several key ways. First, the externality in our model arises from information leakage and privacy costs rather than inventory or production risk, leading to a fundamentally different welfare trade-off. Second, while classical models feature linear or convex demand and cost functions, the data buyer’s valuation in our model is submodular across platforms, capturing diminishing marginal returns to heterogeneous data sources. Finally, instead of price or quantity coordination, the platforms strategically choose their information quality (noise levels) and prices, mediating user participation and buyer demand.
@@
The rest of the paper proceeds as follows. In Section 2, we describe our model and the interactions among the platforms, the data buyer, and the user. We also present some properties of the revealed information measure in our setting, such as its monotonicity and submodularity. In Section 3, we introduce our equilibrium concept and provide preliminary characterizations of it. In Section 4, we focus on a setting with two platforms and provide a full characterization of the equilibrium. The focus on two platforms allows us to describe some of the nuances of the interactions in our model. In Section 5, we establish that our main characterizations continue to hold in the general setting with any number of platforms. We also provide several insights and comparative statics, leading to our discussion of the regulations in Section 6. Finally, Section 7 concludes while the appendix contains all the omitted proofs from the text.
@@ 2 Model
We consider $K\in\mathbb{N}$ platforms that are interacting with a user and a data buyer. The user benefits from using these platforms’ services. In return, the platforms acquire data about the user’s preferences and behavior. The platforms may subsequently sell this data to a third party, such as an advertiser, which we refer to as a data buyer (or, interchangeably, a buyer). The data buyer is interested in learning about the user’s preferences, which can adversely impact the user’s utility. This negative impact can be because of price discrimination, unfairly targeted advertising, manipulation, or intrinsic privacy losses. In the example of MasterCard described in the introduction, MasterCard is one of the platforms, and the data buyer is any other firm that purchases user data from MasterCard.
Formally, we represent the user’s data by $\bm{\theta}\in\mathbb{R}^{d}$, which follows a Gaussian distribution with zero mean and identity variance; that is, $\bm{\theta}\sim\mathcal{N}(0,\mathbb{I}_{d})$. This vector $\bm{\theta}$ can be interpreted as the user’s feature vector. Each platform $i\in[K]\triangleq\{1,\dots,K\}$ offers a service whose quality depends on how accurately the platform can learn the match $\bm{x}_{i}^{\top}\bm{\theta}$ between the user’s feature vector and the platform’s characteristics. Here, $\bm{x}_{i}\in\mathbb{R}^{d}$ is a known unit-norm vector representing the characteristics of platform $i$.
@@
The noise that this signal introduces ensures a degree of privacy with respect to the user’s data. [^1] We set the noise variance to one to simplify the results. The analysis remains valid for any (potentially heterogeneous) level of noise.
The data buyer aims to estimate a linear function of the user’s data, $\bm{y}^{\top}\bm{\theta}$. Here, $\bm{y}\in\mathbb{R}^{d}$ with $\|\bm{y}\|=1$ denotes the data buyer’s known characteristic vector, where here and throughout the paper we use the two-norm. If platform $i$ possesses the user’s data (i.e., $a_{i}=1$), they present the data buyer with an option to acquire a noisy version of their data at a price $p_{i}$. Therefore, the data buyer strategy profile is $\bm{b}\in\{0,1\}^{K}$, with $b_{i}=1$ denoting their acceptance of the platform $i$’s offer. The data buyer then receives a signal
@@
for the price $p_{i}$, where $\sigma_{i}^{2}$ represents the variance of the noise that platform $i$ adds to the data they share with the data buyer. In light of the discussion on differential privacy in the introduction, adding Gaussian noise—known as the Gaussian mechanism—makes the user’s data locally differentially private. This means that a change in a user’s data has a limited impact on the distribution of the shared signal $\tilde{s}_{i}$, where the extent of the impact is controlled by the variance $\sigma_{i}^{2}$. This, in turn, limits how much information the signal $\tilde{s}_{i}$ reveals about the user’s data to the data buyer. [^2] More formally, a randomized mapping $\mathcal{M}(\cdot)$ is said to be $(\varepsilon,\delta)$-locally differentially private if $\mathbb{P}(\mathcal{M}(s)\in W)\leq\exp(\varepsilon)\mathbb{P}(\mathcal{M}(s^{\prime})\in W)+\delta,$ for any $s,s^{\prime}$ and any measurable set $W$. See Duchi [2016] for further details and the characterization of the Gaussian mechanism.
@@
We set $\tilde{s}_{i}=\text{NA}$ when the data buyer does not obtain any information from platform $i$. This can occur either because the user has not shared their data with platform $i$ or because the data buyer declines the offer from platform $i$ or the platform has not provided the service to the user. In this case, $p_{i}$ is irrelevant, and we set it to zero. A summary of the model is illustrated in Fig. 1.
Figure 1: A Summary of the Model.
@@
denote the correlation between platform $i$’s data and the data buyer. Then, we assume
@@
Note that, under this assumption, the data offered by different platforms are viewed as substitutes for each other from the data buyer’s perspective. As an example, suppose the data buyer is interested in inferring whether a household has children. The buyer collects data from both a streaming platform and an online retailer. The overlapping component of these two data sources such as evidence of child-oriented content consumption or purchases is precisely the information about whether the family has children. Once this common factor is taken into account, the remaining parts of the two data streams are largely unrelated, making the datasets substitutes rather than complements.
However, if this assumption does not hold, a complementary effect could also be observed. To illustrate this, suppose we have $K=2$ platforms where $x_{1}$ is orthogonal to $y$, i.e., $\gamma_{1}=0$, and also $x_{2}=\gamma_{2}y+(1-\gamma_{2})x_{1}$. In this scenario, the data from platform 1 alone has no value to the data buyer, as $x_{1}^{\top}\theta$ provides no information about $\theta^{\top}y$. However, if the data buyer decides to purchase data from the second platform, then the data from the first platform could also become valuable. To understand why, observe that the data from the second platform is a convex combination of two components: $x_{1}^{\top}\theta$ and $y^{\top}\theta$. Therefore, if the data buyer acquires the data from the second platform, learning $x_{1}^{\top}\theta$ could assist in better distinguishing these two components. In other words, the data from the first platform is valuable only if accompanied by the data from the second platform. This scenario exemplifies the complementary effect we described earlier. Assumption 1 excludes this possibility, allowing us to primarily focus on the substitution effect. This assumption also implies that the correlation among data for different platforms is given by
$\bm{x}_{i}^{\top}\bm{x}_{j}=\gamma_{i}\gamma_{j}.$